# Dents in the Mirror: A Novel Probe of Dark Matter Substructure in Galaxy Clusters from the Astrometric Asymmetry of Lensed Arcs

### v1

@@ 3.2.1 Subhalo Mass Function
Figure 4: The infall (dashed) and bound (solid) subhalo mass functions that we sample from at each realization for our simulation. For illustrative purposes, darker shades of blue correspond to increasing $\Sigma_{\rm sub}$. The infall SHMF is restricted to subhalos in the mass range $6<\log(m/M_{\odot})<10$. The bound SHMF is calculated from the tidal stripping model presented in du25. Shaded regions indicate the $1\sigma$ scatter in sampling the SHMF.
We use the open-souce software pyHalo [^3] https://github.com/dangilman/pyHalo (gilman20) to populate the region near lensed images with cluster subhalos distributed with a constant surface mass density near a critical curve. We draw cluster subhalo masses from the following SHMF:
@@
where $\Sigma_{\rm sub}$ is the SHMF normalization (in units of kpc-2 throughout unless otherwise stated), $\alpha=1.9$ (Springel2008; giocoli08), and the pivot mass $m_{0}=10^{8}M_{\odot}$. The SHMF represents the infall SHMF (thus $m$ is equal to the infall subhalo mass), which is the mass function of accreted halos prior to tidal evolution. The infall SHMF is predicted by CDM to be a universal property for subhalos (giocoli08), whereas the evolved SHMF (or equivalently the bound SHMF) depends on the cluster host properties and subhalo infall trajectories (han16). Figure 4 shows the SHMF for various $\Sigma_{\rm sub}$. We only sample subhalos between the infall mass range $6<\log(m/M_{\odot})<10$, consistent with previous works at cluster scales (dai20a; williams24). Subhalos above this range are rare, and likely would host a visible galaxy that we would identify as a cluster member. Below this range, halos produce negligible contributions to the deflection field (dai18). Given the low redshifts of the clusters we consider, most perturbers are expected to be subhalos, and we leave the inclusion of line-of-sight halos to future work.
After infall, subhalos lose mass to tidal stripping, and retain a fraction $f_{\rm{bound}}$ of their infall mass, with a final mass $m_{\rm{bound}}=f_{\rm{bound}}m$. In principle, a subhalo population will have a distribution of $f_{\rm{bound}}$, depending on individual subhalo orbits, concentrations, and infall times (han16; du25). Predicting the distribution of $f_{\rm{bound}}$ for a given cluster requires developing a tidal evolution model for cluster subhalos to account for these factors, which could follow the approach used for galaxy-scale lens systems (du25).
Lacking a detailed tidal evolution model for cluster subhalos, we define $\bar{f}_{\rm bound}$ as the mean bound mass fraction for cluster subhalos. In our inference procedure, we treat the mean bound mass fraction $\bar{f}_{\rm bound}$ as a free parameter, alongside the normalization of the infall mass function $\Sigma_{\rm{sub}}$. We set the mean bound mass fraction with a log-uniform prior $\log\left(\bar{f}_{\rm bound}\right)\sim\mathcal{U}\left(-1.25,-0.75\right)$, comparable to the distribution of bound masses for a group-scale lens (du25). For each realization, the subhalos will have a normally distributed $f_{\rm bound}$, with mean $\bar{f}_{\rm bound}$ and standard deviation of each log-distribution to be 0.5. This corresponds to a mean bound mass fraction range between $\sim 5-20\%$. For $\Sigma_{\rm sub}$, we set a wide log-uniform prior $\log\left(\Sigma_{\rm sub}\right)\sim\mathcal{U}\left(-3.5,-0.1\right)$.
@@ 3.2.2 Subhalo Density Profiles
Figure 5: Summary plots of the tidal stripping process and how it manifests in the subhalo density profiles. For this example, the subhalo population is simulated to have a Gaussian bound mass fraction distribution: $\log f_{\rm bound}\sim\mathcal{N}\left(-1.0,0.5\right)$. We note that this mean bound mass fraction is the center of the log-uniform prior we set for $\bar{f}_{\rm bound}$. From this distribution, the tidal evolution tracks from du25 are used to calculate the each subhalo’s corresponding truncation radius $r_{t}$, and normalization $f_{t}$, with some scatter. These tidal tracks can be visualized as a tight relation between $f_{\rm bound}$ and $r_{t}$ (Top) and $f_{\rm bound}$ and $f_{t}$ (Bottom). In both panels, blue contours are logarithmically spaced 2D density distributions for the parameters, red dots are an example subhalo population (for illustrative purposes $\log\Sigma_{\rm sub}=-1.0$ is shown, although the tidal tracks are the same for any value of $\Sigma_{\rm sub}$), and dashed lines are the means of the parameter distributions (top and right histograms).
Once the bound mass is known, and normally distributed according to the defined $f_{\rm bound}$ distribution described in Section 3.2.1, the tidal tracks (errani21; du24) are used to calculate each subhalo’s density profile. We simulate each subhalo as a tidally stripped NFW. For CDM, we model this as a truncated NFW:
@@
where $\rho_{s}$ is the scale density, $r_{s}$ is the scale radius, and $f_{t}$ and $r_{t}$ are truncation parameters determined by the bound mass fraction $f_{\rm bound}$ after tidal stripping (du24).
The effect of the tidal stripping process manifests in the density profile as an effective rescaling and truncation. The tidal tracks that govern this can be found in du24, where our parametrization of the subhalo density profile matches their Nuker model with $\alpha=1$, $\beta=3$, $\gamma=1$, and $\delta=2$ (see equation 8 of du24). As in du25, we do not employ the exact relations between $f_{\rm bound}$, $f_{t}$, and $r_{t}$, due to redshift dependency of the halo mass. Instead, following du25, we first convert $f_{\rm bound}$ to $f_{\rm bound,mx}\equiv m_{\rm bound}/m_{\rm mx,0}$, and then adopt the relation between $f_{\rm bound,mx}$, $f_{t}$ and $r_{t}$. This relation has been shown to be more universal and is independent of the virial mass definition across different redshifts. Here, $m_{\rm mx,0}$ is the mass enclosed within $r_{\rm max,0}$ at the time of infall, where $r_{\rm max,0}$ is the radius at which the circular velocity $v_{c}$ reaches its maximum value:
@@
In this procedure, $r_{s}$ and $\rho_{s}$ are calculated with respect to $M_{200}$ at the critical density of the Universe at the infall redshift. We derive the distribution of infall redshifts for a $10^{14}M_{\odot}$ host halo using galacticus (benson12). The universal tidal track to compute the corresponding $m_{\rm bound}$ is obtained from the mass enclosed within the radius where the maximum $v_{c}$ is reached. This allows us to directly calculate the specified $f_{\rm bound}$ distribution from a universal tidal track at any infall redshift. The resulting relations between $f_{\rm bound}$, $f_{t}$, and $r_{t}$ that we use in this work are shown in Figure 5.
@@ 3.4 Bayesian Inference of the Cluster Mass Fraction
where $A_{\rm ap}$ is the projected area of the lens plane that we are simulating dark matter subhalos (i.e. the area of the simulation aperture). $M_{\rm low}$ and $M_{\rm high}$ are the low and high subhalo mass limits that we sample the SHMF between, respectively. As mentioned in Section 3.2.1, these are set to be $\log(M_{\rm low}/M_{\odot})=6$ and $\log(M_{\rm high}/M_{\odot})=10$. It is worth noting that $f_{\rm sub}$ is calculated directly by integrating each sampled SHMF. We fix $\alpha=1.9$ for this analysis, as it is a robust prediction from N-body simulations in CDM (giocoli08; Springel2008). $f_{\rm sub}$ can be analytically written as:
@@
Even in this framework, however, the use of ABC is still intractable. This is primarily due to the modeling of the macrolens parameters $\boldsymbol{M}$. In principle, at each sampling iteration a different set of $\boldsymbol{M}$ would be generated from a lens model fit, as has been done in similar analyses at galaxy scales (gilman20). On cluster scales, this is computationally intractable, as lens models for clusters often take multiple days to run. We therefore proceed with the approximation of fixing the background macrolens during the ABC process. This turns out to a be a reasonable approximation for this problem since the variations in the macrolens will not affect the asymmetry of an arc, as quantified by $\xi$.
The idea here is to reduce the dimensionality of the problem and to ensure that we are adequately capturing the relevant information from each realization, which in our case is the asymmetry (equation 22). Thus, we replace the formal likelihood function with an approximation based on $w\left(\rho\right)$, which depends on the model parameters $\boldsymbol{m}$ and $\boldsymbol{M}$ through the model-predicted datasets $\boldsymbol{D^{\prime}}$. In our analysis procedure, we set $\epsilon$ implicitly by generating a large number $N$ of model-predicted datasets $\boldsymbol{D^{\prime}}$, and accepting the top 100 samples. Our ABC analysis proceeds as follows (in order):
@@
  • The preceding steps are repeated for $N$ realizations. We obtain an approximation of the posterior distribution by accepting the 100 samples corresponding to the lowest distance metrics $\rho$.
@@ 3.5 Joint Constraints from Multiple Galaxy Clusters
We can use the self-similarity of halo substructure in CDM to generalize our model to a cluster of any virial mass and density profile. Following han16, we can write
@@
where $R$ is the projected distance from the cluster mass center where we observed lensed images, $R_{s}$ is the scale radius of the host, and $\gamma\sim 1$ is a parameter that encodes the amount of tidal stripping experienced by subhalos with different projected distances to the cluster center. The function $\tilde{\kappa}_{\rm{host}}\left(M_{\rm{host}},R/R_{s}\right)$ varies in proportion to the projected mass density of the host halo, and captures the radial and cluster-mass dependence of the projected surface mass density of subhalos. The term $\left(R/R_{s}\right)^{\gamma}$ encodes the radial dependence of the mean bound mass fraction. Note that tidal stripping of CDM subhalos appears approximately independent of the subhalo infall mass (e.g. du25).
@@
This approach lays the groundwork for future inferences of the cluster subhalo mass function from multiple systems with variations in $M_{\rm{host}}$ and $R_{s}$ among the sample. In this first analysis, we will apply the methodology to two cluster arcs, AS1063, and MACSJ0416, and derive an inference of $f_{\rm{sub}}$. If the total mass in dark matter $M_{\rm{ap}}$ (see Equation 24) varies in proportion to $\tilde{\kappa}\left(M_{\rm{host}},R/R_{s}\right)$, and assuming dark matter dominates the surface mass density near the cluster critical curve, the subhalo mass fraction between the different arcs will differ only by the factor $\left(R/R_{s}\right)^{\gamma}$. When combining inferences of $f_{\rm{sub}}$, we assume that we are measuring the mass fraction at approximately the same $R/R_{s}$, such that we can directly multiply $f_{\rm{sub}}$ likelihoods obtained from each arc. This approximation is sufficient to first order, as estimates from CLASH measure $M_{\rm 200}$ and $R_{\rm 200}$ for both clusters to be $\sim 10^{15}M_{\odot}$ and $\sim 2$ Mpc, respectively (umetsu14). Future analyses will account for the scaling with cluster mass and radius described by Equation 28.
@@ 5.1 Data
Figure 10: The two arcs we consider in this work: AS1063 System 1 (Left Column) and the MACSJ0416 Warhol Arc (Right Column). The top row shows images for both arcs, with an HST F606W image for AS1063 System 1 (lotz17) and a JWST F090W image (windhorst23) for the Warhol arc. In each case, North is up and East is left. Open circles indicate the observed image positions, with knots sharing color representing counterimaged pairs. The bottom row shows the modeling window used for our analysis for each arc. The window depicted is the same for each arc as their image in the top row. For AS1063 System 1 and the Warhol Arc, we use the lens models from bergamini19 and rihtarsic25 as the macrolens density, respectively. The blue squares show the observed image positions, which correspond to the open circles in the image in the top row. Green diamonds and yellow triangles indicate the reconstructed images from the lens model and predicted midpoints, respectively. The light red dashed line is the macrolens critical curve. Purple contours show the logarithmically spaced contours of the macrolens density profile. The $x$ and $y$ positions are presented in arcseconds with respect to RA,Dec = (342.1832, -44.5309) and (64.0382, -24.0675) for AS1063 System 1 and the Warhol Arc, respectively.
@@
We consider two arcs in this work: Abell S1063 (hereafter AS1063) System 1 and the MACS J0416.1-2403 (hereafter MACSJ0416) Warhol Arc. AS1063 and MACSJ0416 have lens redshifts $z_{d}$ of 0.348 (guzzo09) and 0.396 (postman12), respectively. The spectroscopically confirmed source redshifts $z_{s}$ for AS1063 System 1 and the Warhol arc are 1.229 (balestra13) and 0.9397 (caminha17), respectively. Both clusters are well studied and were extensively observed with the Hubble Frontier Fields (HFF) program (lotz17). Figure 10 shows an HST image of AS1063 System 1 and a JWST image of the MACSJ0416 Warhol Arc.
@@
For AS1063, there have been numerous lens models published in the last decade (diego16b; bergamini19; limousin22). For this work, we adopt the model from bergamini19 (hereafter B19) to act as the macrolens for this analysis. The B19 lens model utilizes the parametric lens inversion algorithm LensTool (kneib96; jullo07), which models the density with smooth elliptical potentials for cluster-scale halos and cluster member galaxies, making it suitable for this analysis. B19 achieves a lens plane RMS of 0.55" with HFF data.
@@
AS1063 System 1 is a large perpendicular arc that spans $\sim 0.3$" along the critical curve, and $\sim 7$" perpendicularly from its critical curve. There are three bright counterimaged knots that we use to measure the observed asymmetry. In previous lensing analyses, however, only a single pair of counterimages has been used to constrain the lens model. Therefore, we measure the image positions of the knots ourselves by simply taking them to be the brightest pixel position of each knot’s flux centroid. We note that this is not the most rigorous procedure to measure the image positions; however, it serves as a reasonable initial approximation for this analysis. We also emphasize that this is roughly consistent with the typical image identification procedure that occurs in lens modelling studies, where new image candidates are identified by color and morphology, then confirmed by their consistency in being reproduced by existing lens models (lotz17). Therefore, we implement this simple procedure as a sufficient approximation, and reserve more sophisticated follow-up image identification procedures for lensed arcs to future work. With the observed image positions (presented in Figure 10 and Table 3), we calculate the source positions by backprojecting the observed image positions with the B19 lens model and taking the mean of the two counterimages to be the model source position. This is then forward projected with the lens model to calculate the reconstructed image position. The reconstructed image positions are the images that we then apply our analysis on (i.e. to generate many realizations of $\hat{\xi}$ to sample the posterior with Equation 26).
@@
For MACSJ0416, the procedure is largely the same as for AS1063. MACSJ0416 has a variety of recent lens models, some of which have made use of recent JWST imaging allowing for >400>400 lensed image constraints (e.g. bergamini23; cha23; diego23a; perera24b; rihtarsic25; limousin25). We adopt the model from rihtarsic25 (hereafter R25) for the macrolens for this arc. R25 makes use of JWST imaging from the CAnadian NIRISS Unbiased Cluster Survey (CANUCS) along with pre-JWST multiple image catalogues (richard21; bergamini23) and also utilizes LensTool for a lens plane RMS of 0.52".
@@
The Warhol Arc (kaurov19; chen19) in MACSJ0416 is a large parallel arc that extends $\sim 2.3$" along the critical curve. It has been the subject of considerable interest in the search for lensed transients, and thus has had numerous follow-up observations conducted on it (kelly22; yan23). Recently, palencia25 examined the effect of compact dark matter objects on the spatial distribution of transient events, establishing the Warhol Arc as a strong candidate to use to study the nature of dark matter. R25 identifies six lensed knots that they use as constraints in their model. For our analysis, we include an additional two lensed knots, which are clearly visible in recent JWST imaging (see Figure 2 of yan23). This gives us a total of eight imaged knots to use for our analysis. Like with AS1063 System 1, we measure each image’s position ourselves by taking them as the brightest pixel position. The same procedure is then repeated to calculate the R25 reconstructed image positions for the eight knots. The image positions for the Warhol Arc are shown in Figure 10 and Table 4.

### v2

@@ 3.2.1 Subhalo Mass Function
Figure 4: The infall (dashed) and bound (solid) subhalo mass functions that we sample from at each realization for our simulation. For illustrative purposes, darker shades of blue correspond to increasing $\Sigma_{\rm sub}$. The infall SHMF is restricted to subhalos in the mass range $6<\log(m/M_{\odot})<10$. The bound SHMF is calculated from the tidal stripping model presented in Du et al. (2025). Shaded regions indicate the $1\sigma$ scatter in sampling the SHMF.
We use the open-souce software pyHalo [^4] https://github.com/dangilman/pyHalo (Gilman et al., 2020) to populate the region near lensed images with cluster subhalos distributed with a constant surface mass density near a critical curve. We draw cluster subhalo masses from the following SHMF:
@@
where $\Sigma_{\rm sub}$ is the SHMF normalization (in units of kpc-2 throughout unless otherwise stated), $\alpha=1.9$ (Springel et al., 2008; Giocoli et al., 2008), and the pivot mass $m_{0}=10^{8}M_{\odot}$. The SHMF represents the infall SHMF (thus $m$ is equal to the infall subhalo mass), which is the mass function of accreted halos prior to tidal evolution. The infall SHMF is predicted by CDM to be a universal property for subhalos (Giocoli et al., 2008), whereas the evolved SHMF (or equivalently the bound SHMF) depends on the cluster host properties and subhalo infall trajectories (Han et al., 2016). Figure 4 shows the SHMF for various $\Sigma_{\rm sub}$. We only sample subhalos between the infall mass range $6<\log(m/M_{\odot})<10$, consistent with previous works at cluster scales (Dai et al., 2020; Williams et al., 2024b). Subhalos above this range are rare, and likely would host a visible galaxy that we would identify as a cluster member. Below this range, halos produce negligible contributions to the deflection field (Dai et al., 2018). We examine how changes in this mass range would affect our results in Appendix C. Given the low redshifts of the clusters we consider, most perturbers are expected to be subhalos, and we leave the inclusion of line-of-sight halos to future work.
After infall, subhalos lose mass to tidal stripping, and retain a fraction $f_{\rm{bound}}$ of their infall mass, with a final mass $m_{\rm{bound}}=f_{\rm{bound}}m$. In principle, a subhalo population will have a distribution of $f_{\rm{bound}}$, depending on individual subhalo orbits, concentrations, and infall times (Han et al., 2016; Du et al., 2025). Predicting the distribution of $f_{\rm{bound}}$ for a given cluster requires developing a tidal evolution model for cluster subhalos to account for these factors, which could follow the approach used for galaxy-scale lens systems (Du et al., 2025).
Lacking a detailed tidal evolution model for cluster subhalos, we define $\bar{f}_{\rm bound}$ as the mean bound mass fraction for cluster subhalos. In our inference procedure, we treat the mean bound mass fraction $\bar{f}_{\rm bound}$ as a free parameter, alongside the normalization of the infall mass function $\Sigma_{\rm{sub}}$. We set the mean bound mass fraction with a log-uniform prior $\log\left(\bar{f}_{\rm bound}\right)\sim\mathcal{U}\left(-1.25,-0.75\right)$, comparable to the distribution of bound masses for a group-scale lens (Du et al., 2025). For each realization, the subhalos will have a normally distributed $f_{\rm bound}$, with mean $\bar{f}_{\rm bound}$ and standard deviation of each log-distribution to be 0.5. This corresponds to a mean bound mass fraction range between $\sim 5-20\%$. For $\Sigma_{\rm sub}$, we set a wide log-uniform prior $\log\left(\Sigma_{\rm sub}\right)\sim\mathcal{U}\left(-3.5,-0.1\right)$.
@@ 3.2.2 Subhalo Density Profiles
Figure 5: Summary plots of the tidal stripping process and how it manifests in the subhalo density profiles. For this example, the subhalo population is simulated to have a Gaussian bound mass fraction distribution: $\log f_{\rm bound}\sim\mathcal{N}\left(-1.0,0.5\right)$. We note that this mean bound mass fraction is the center of the log-uniform prior we set for $\bar{f}_{\rm bound}$. From this distribution, the tidal evolution tracks from Du et al. (2025) are used to calculate the each subhalo’s corresponding truncation radius $r_{t}$, and normalization $f_{t}$, with some scatter. These tidal tracks can be visualized as a tight relation between $f_{\rm bound}$ and $r_{t}$ (Top) and $f_{\rm bound}$ and $f_{t}$ (Bottom). In both panels, blue contours are logarithmically spaced 2D density distributions for the parameters, red dots are an example subhalo population (for illustrative purposes $\log\Sigma_{\rm sub}=-1.0$ is shown, although the tidal tracks are the same for any value of $\Sigma_{\rm sub}$), and dashed lines are the means of the parameter distributions (top and right histograms).
Once the bound mass is known, and normally distributed according to the defined $f_{\rm bound}$ distribution described in Section 3.2.1, the tidal tracks (Errani and Navarro, 2021; Du et al., 2024) are used to calculate each subhalo’s density profile. We simulate each subhalo as a tidally stripped NFW. For CDM, we model this as a truncated NFW:
@@
where $\rho_{s}$ is the scale density, $r_{s}$ is the scale radius, and $f_{t}$ and $r_{t}$ are truncation parameters determined by the bound mass fraction $f_{\rm bound}$ after tidal stripping (Du et al., 2024).
The effect of the tidal stripping process manifests in the density profile as an effective rescaling and truncation. The tidal tracks that govern this can be found in Du et al. (2024), where our parametrization of the subhalo density profile matches their Nuker model with $\alpha=1$, $\beta=3$, $\gamma=1$, and $\delta=2$ (see equation 8 of Du et al. (2024)). As in Du et al. (2025), we do not employ the exact relations between $f_{\rm bound}$, $f_{t}$, and $r_{t}$, due to redshift dependency of the halo mass. Instead, following Du et al. (2025), we first convert $f_{\rm bound}$ to $f_{\rm bound,mx}\equiv m_{\rm bound}/m_{\rm mx,0}$, and then adopt the relation between $f_{\rm bound,mx}$, $f_{t}$ and $r_{t}$. This relation has been shown to be more universal and is independent of the virial mass definition across different redshifts. Here, $m_{\rm mx,0}$ is the mass enclosed within $r_{\rm max,0}$ at the time of infall, where $r_{\rm max,0}$ is the radius at which the circular velocity $v_{c}$ reaches its maximum value:
@@
In this procedure, $r_{s}$ and $\rho_{s}$ are calculated with respect to $M_{200}$ at the critical density of the Universe at the infall redshift. We derive the distribution of infall redshifts for a $10^{14}M_{\odot}$ host halo using galacticus (Benson, 2012). The universal tidal track to compute the corresponding $m_{\rm bound}$ is obtained from the mass enclosed within the radius where the maximum $v_{c}$ is reached. This allows us to directly calculate the specified $f_{\rm bound}$ distribution from a universal tidal track at any infall redshift. The resulting relations between $f_{\rm bound}$, $f_{t}$, and $r_{t}$ that we use in this work are shown in Figure 5. We note that the distribution of 3D radii of subhalos near the critical curve affects the distribution of the bound mass fraction for a given subhalo population. This projection effect is accounted for in multiple ways. We account for it in our modeling framework by setting the mean bound mass fraction as a free parameter which we marginalize over. It is also accounted for when calculating the infall redshift distribution.
@@ 3.4 Bayesian Inference of the Cluster Mass Fraction
where $A_{\rm ap}$ is the projected area of the lens plane that we are simulating dark matter subhalos (i.e. the area of the simulation aperture). $M_{\rm low}$ and $M_{\rm high}$ are the low and high subhalo mass limits that we sample the SHMF between, respectively. As mentioned in Section 3.2.1, these are set to be $\log(M_{\rm low}/M_{\odot})=6$ and $\log(M_{\rm high}/M_{\odot})=10$. It is worth noting that $f_{\rm sub}$ is calculated directly by integrating each sampled SHMF. We fix $\alpha=1.9$ for this analysis, as it is a robust prediction from N-body simulations in CDM (Giocoli et al., 2008; Springel et al., 2008). $f_{\rm sub}$ can be analytically written as:
@@
Even in this framework, however, the use of ABC is still intractable. This is primarily due to the modeling of the macrolens parameters $\boldsymbol{M}$. In principle, at each sampling iteration a different set of $\boldsymbol{M}$ would be generated from a lens model fit, as has been done in similar analyses at galaxy scales (Gilman et al., 2020). On cluster scales, this is computationally intractable, as lens models for clusters often take multiple days to run. We therefore proceed with the approximation of fixing the background macrolens during the ABC process. This turns out to a be a reasonable approximation for this problem since the variations in the macrolens will not affect the asymmetry of an arc, as quantified by $\xi$.
The idea here is to reduce the dimensionality of the problem and to ensure that we are adequately capturing the relevant information from each realization, which in our case is the asymmetry (equation 22). Thus, we replace the formal likelihood function with an approximation based on $w\left(\rho\right)$, which depends on the model parameters $\boldsymbol{m}$ and $\boldsymbol{M}$ through the model-predicted datasets $\boldsymbol{D^{\prime}}$. In our analysis procedure, we set $\epsilon$ implicitly by generating a large number $N$ of model-predicted datasets $\boldsymbol{D^{\prime}}$, and accepting the top 100 samples. We verify that our ABC analysis yields converged posterior distributions in Appendix D. Our ABC analysis proceeds as follows (in order):
@@
  • The preceding steps are repeated for $N$ realizations, taken to be 20000 in this work. We obtain an approximation of the posterior distribution by accepting the 100 samples corresponding to the lowest distance metrics $\rho$ (a tolerance threshold of 0.5%). We emphasize that this statistical framework accounts for the relative likelihood that $\xi$ is dominated by deflection from a single massive subhalo by averaging over many realizations. We explicitly examine this effect in Appendix C.
@@ 3.5 Joint Constraints from Multiple Galaxy Clusters
We can use the self-similarity of halo substructure in CDM to generalize our model to a cluster of any virial mass and density profile. Following Han et al. (2016), we can write
@@
where $R$ is the projected distance from the cluster mass center where we observed lensed images, $R_{s}$ is the scale radius of the host, and $\gamma\sim 1$ is a parameter that encodes the amount of tidal stripping experienced by subhalos with different projected distances to the cluster center. The function $\tilde{\kappa}_{\rm{host}}\left(M_{\rm{host}},R/R_{s}\right)$ varies in proportion to the projected mass density of the host halo, and captures the radial and cluster-mass dependence of the projected surface mass density of subhalos. The term $\left(R/R_{s}\right)^{\gamma}$ encodes the radial dependence of the mean bound mass fraction. Note that tidal stripping of CDM subhalos appears approximately independent of the subhalo infall mass (e.g. Du et al., 2025).
@@
This approach lays the groundwork for future inferences of the cluster subhalo mass function from multiple systems with variations in $M_{\rm{host}}$ and $R_{s}$ among the sample. In this first analysis, we will apply the methodology to two cluster arcs, AS1063, and MACSJ0416, and derive an inference of $f_{\rm{sub}}$. If the total mass in dark matter $M_{\rm{ap}}$ (see Equation 24) varies in proportion to $\tilde{\kappa}\left(M_{\rm{host}},R/R_{s}\right)$, and assuming dark matter dominates the surface mass density near the cluster critical curve, the subhalo mass fraction between the different arcs will differ only by the factor $\left(R/R_{s}\right)^{\gamma}$. When combining inferences of $f_{\rm{sub}}$, we assume that we are measuring the mass fraction at approximately the same $R/R_{s}$, such that we can directly multiply $f_{\rm{sub}}$ likelihoods obtained from each arc. This approximation is sufficient to first order, as estimates from CLASH measure $M_{\rm 200}$ and $R_{\rm 200}$ for both clusters to be $\sim 10^{15}M_{\odot}$ and $\sim 2$ Mpc, respectively (Umetsu et al., 2014). Future analyses will account for the scaling with cluster mass and radius described by Equation 28.
@@ 5.1 Data
Figure 10: The two arcs we consider in this work: AS1063 System 1 (Left Column) and the MACSJ0416 Warhol Arc (Right Column). The top row shows images for both arcs, with an HST F606W image for AS1063 System 1 (Lotz et al., 2017) and a JWST F090W image (Windhorst et al., 2023) for the Warhol arc. In each case, North is up and East is left. Open circles indicate the observed image positions, with knots sharing color representing counterimaged pairs. The bottom row shows the modeling window used for our analysis for each arc. The window depicted is the same for each arc as their image in the top row. For AS1063 System 1 and the Warhol Arc, we use the lens models from Bergamini et al. (2019) and Rihtaršič et al. (2025) as the macrolens density, respectively. The blue squares show the observed image positions, which correspond to the open circles in the image in the top row. Green diamonds and yellow triangles indicate the reconstructed images from the lens model and predicted midpoints, respectively. The light red dashed line is the macrolens critical curve. Purple contours show the logarithmically spaced contours of the macrolens density profile. The $x$ and $y$ positions are presented in arcseconds with respect to RA,Dec = (342.1832, -44.5309) and (64.0382, -24.0675) for AS1063 System 1 and the Warhol Arc, respectively.
@@
We consider two arcs in this work: Abell S1063 (hereafter AS1063) System 1 and the MACS J0416.1-2403 (hereafter MACSJ0416) Warhol Arc. AS1063 and MACSJ0416 have lens redshifts $z_{d}$ of 0.348 (Guzzo et al., 2009) and 0.396 (Postman et al., 2012), respectively. The spectroscopically confirmed source redshifts $z_{s}$ for AS1063 System 1 and the Warhol arc are 1.229 (Balestra et al., 2013) and 0.9397 (Caminha et al., 2017), respectively. Both clusters are well studied and were extensively observed with the Hubble Frontier Fields (HFF) program (Lotz et al., 2017). Figure 10 shows an HST image of AS1063 System 1 and a JWST image of the MACSJ0416 Warhol Arc.
@@
For AS1063, there have been numerous lens models published in the last decade (Diego et al., 2016; Bergamini et al., 2019; Limousin et al., 2022). For this work, we adopt the model from Bergamini et al. (2019) (hereafter B19) to act as the macrolens for this analysis. The B19 lens model utilizes the parametric lens inversion algorithm LensTool (Kneib et al., 1996; Jullo et al., 2007), which models the density with smooth elliptical potentials for cluster-scale halos and cluster member galaxies, making it suitable for this analysis. B19 achieves a lens plane RMS of 0.55" with HFF data.
@@
AS1063 System 1 is a large perpendicular arc that spans $\sim 0.3$" along the critical curve, and $\sim 7$" perpendicularly from its critical curve. There are three bright counterimaged knots that we use to measure the observed asymmetry. In previous lensing analyses, however, only a single pair of counterimages has been used to constrain the lens model. Therefore, we measure the image positions of the knots ourselves by simply taking them to be the brightest pixel position of each knot’s flux centroid. We note that this is not the most rigorous procedure to measure the image positions; however, it serves as a reasonable initial approximation for this analysis. We also emphasize that this is roughly consistent with the typical image identification procedure that occurs in lens modelling studies, where new image candidates are identified by color and morphology, then confirmed by their consistency in being reproduced by existing lens models (Lotz et al., 2017). Therefore, we implement this simple procedure as a sufficient approximation, and reserve more sophisticated follow-up image identification procedures for lensed arcs to future work. With the observed image positions (presented in Figure 10 and Table 3), we calculate the source positions by backprojecting the observed image positions with the B19 lens model and taking the mean of the two counterimages to be the model source position. This is then forward projected with the lens model to calculate the reconstructed image position. The reconstructed image positions are the images that we then apply our analysis on (i.e. to generate many realizations of $\hat{\xi}$ to sample the posterior with Equation 26).
@@
For MACSJ0416, the procedure is largely the same as for AS1063. MACSJ0416 has a variety of recent lens models, some of which have made use of recent JWST imaging allowing for >400>400 lensed image constraints (e.g. Bergamini et al., 2023; Cha and Jee, 2023; Diego et al., 2024; Perera et al., 2025b; Rihtaršič et al., 2025; Limousin et al., 2025). We adopt the model from Rihtaršič et al. (2025) (hereafter R25) for the macrolens for this arc. R25 makes use of JWST imaging from the CAnadian NIRISS Unbiased Cluster Survey (CANUCS) along with pre-JWST multiple image catalogues (Richard et al., 2021; Bergamini et al., 2023) and also utilizes LensTool for a lens plane RMS of 0.52".
@@
The Warhol Arc (Kaurov et al., 2019; Chen et al., 2019) in MACSJ0416 is a large parallel arc that extends $\sim 2.3$" along the critical curve. It has been the subject of considerable interest in the search for lensed transients, and thus has had numerous follow-up observations conducted on it (Kelly et al., 2022; Yan et al., 2023). Recently, Palencia et al. (2025) examined the effect of compact dark matter objects on the spatial distribution of transient events, establishing the Warhol Arc as a strong candidate to use to study the nature of dark matter. R25 identifies six lensed knots that they use as constraints in their model. For our analysis, we include an additional two lensed knots, which are clearly visible in recent JWST imaging (see Figure 2 of Yan et al. (2023)). This gives us a total of eight imaged knots to use for our analysis. Like with AS1063 System 1, we measure each image’s position ourselves by taking them as the brightest pixel position. The same procedure is then repeated to calculate the R25 reconstructed image positions for the eight knots. The image positions for the Warhol Arc are shown in Figure 10 and Table 4.